OpenStudy (anonymous):

Find the rectangular coordinates of the point whose polar coordinates are (-3,5 pi/3). a.) (1/2,-sqrt(3)/2) b.) (-1/2, sqrt(3)/2) c.) (-3/2,3 sqrt(3)/2) d.) (3/2,-3 sqrt(3)/2)

OpenStudy (mathmale):

In your shoes, I'd draw this situation. Have you tried doing that?

OpenStudy (anonymous):

i dont understand how to do it.

OpenStudy (mathmale):

I'd suggest you draw a circle of radius 3. Then, draw on this same graph the angle 5pi/3 (300 degrees).

zepdrix (zepdrix):

If you're having trouble drawing it out ( I know that understanding trig conceptually can be difficult ) then you can trust in your identities for converting from polar to Cartesian. \[\Large\bf\sf x=r \cos \theta\]\[\Large\bf\sf y=r \sin \theta\]And we were given the point,\[\Large\bf\sf (r,\theta)\quad=\quad \left(-3,\frac{5\pi}{3}\right)\]

zepdrix (zepdrix):

Just plug the r and theta into the formulas and simplify.

OpenStudy (anonymous):

okay.

zepdrix (zepdrix):

Oh I guess you still have to remember your special angles like cos(5pi/3) :3

OpenStudy (anonymous):

yeah, I got it. Thank you so much.

zepdrix (zepdrix):

Oh good! :)